OpenStudy (anonymous):

Find all solutions of the equation and express them in the form a + bi. x^2 − 8x + 17 = 0 x = ____ (positive imaginary part) x = ____ (negative imaginary part)

OpenStudy (anonymous):

4 + i 4 - i

OpenStudy (anonymous):

--8+sqrt64-68 / 2 = 8 + sqrt-4 /2 = 8 + 2i /2 = 4 + i = positive imag part other root = 4 - i

OpenStudy (anonymous):

can you guys help me with this one? Find all solutions of the equation and express them in the form a + bi. 49x^2 + 3 = 14x x = ____ (positive imaginary part) x = ____ (negative imaginary part)

OpenStudy (anonymous):

quadratic formula gives \[x=\frac{1\pm \sqrt{2}i}{7}\]

OpenStudy (anonymous):

i got that far but whats next

OpenStudy (anonymous):

x = (1 + sqrt2*i)/7 (positive imaginary part) x = (1 - sqrt2*i)/7 (negative imaginary part)

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What makes you doubt satellite73's answer?

OpenStudy (anonymous):

because I believe it is supposed to be further simplified

OpenStudy (anonymous):

i am pretty sure of my answer but it is not in "standard form"

OpenStudy (anonymous):

cannot "simplify" but can write it as \[a+bi\]

OpenStudy (anonymous):

\[\frac{1}{7}+\frac{\sqrt{2}}{7}i\] and its conjugate